OpenStudy (anonymous):

0 = x^3 - x , sorry i just don't even know where to begin on this one

OpenStudy (mimi_x3):

x(x^2-1) x = 0 , 1

OpenStudy (anonymous):

well, your best bet is to factor out an x and get: x(x^2-1)=0

OpenStudy (anonymous):

Wow, that simple? You guys rock

OpenStudy (mimi_x3):

lols , yup its simple xD

OpenStudy (anonymous):

Wait it would be +- 1 correct?

OpenStudy (anonymous):

well, set each factor to zero and solve for x: x(x+1)(x-1) x=0 x+1=0 and x-1=0

OpenStudy (mimi_x3):

(x-1)(x+1) Therefore , its 1, -1

OpenStudy (anonymous):

ok ok awesome, I get it now. Ever since summer my brain has been fried. Thanks guys

OpenStudy (anonymous):

They're wrong, actually. In fact, I think they might be purposefully misleading you. The answer is x = 3.141592654 y = 6.283185307

OpenStudy (mimi_x3):

wth!? your misleading her , how can x be a decimal ?

OpenStudy (anonymous):

The book says the answer is 0,1,-1 soo..............

OpenStudy (anonymous):

read the link and weep

OpenStudy (anonymous):

Because there are an infinite number of values on a coordinate plane and 3.141592654 is one of them

OpenStudy (anonymous):

Wth I didn't know wolframalpha could do that!

OpenStudy (anonymous):

oh man did i just open Pandoras box

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

Wolframalpha is a site operated by a charlatan who's out to deceive school children and lower the national GPA. Don't go to that site.

OpenStudy (anonymous):

warning: use only to check your work, or else your brain may rot

OpenStudy (mimi_x3):

LOL, wth its a gr8 site to get answers straight away

OpenStudy (anonymous):

hahhahah LOL I know thanks

OpenStudy (anonymous):

Find. Don't heed my warning.

OpenStudy (anonymous):

is this supposed to be humor? 3.141592654?

OpenStudy (anonymous):

wow @EbnorEqvine you took the cake on that one

OpenStudy (anonymous):

Hah.